What is the molarity of a bleach solution containing 12.5g of NaOCl per liter of bleach?


1
Expert's answer
2021-03-04T06:33:28-0500

the molecular weight of NaOC l is (22.9+16+34.45)=73.35

(22.9+16+34.45)=73.35this is from adding all the atomic weights from a periodic table. You then divide how many grams by the molecular mass (12.5/73.35)=.1 moles. therefore the molarity is .1/L. For the calcium hydroxide the molecular mass is 40+(2∗(16+1.008))=74.016

40+(2∗(16+1.008))=74.016This is how many grams one mole of calcium hydroxide weighs so .25∗1.5=0.375moles0.375∗74.016

=27.756grams